A rifle bullet, of mass 18.00 g, is fired to the right at a speed of 240.00 m / s to hit a sack of sand and penetrate it 35.00 cm. Determine the magnitude of the friction force (assumed constant) acting on the bullet.

a) –6.17 N

b) 5924.57 N​​​​​​​

c) 1481.14 N​​​​​​​

d) –12.34 N​​​​​​​

e) None of the above.
Answer #1

To determine the constant friction force acting on a bullet as it penetrates a sandbag, we can use the work-energy principle.

Given:

  • Mass of the bullet, m=18.00g=0.018kgm = 18.00 \, \text{g} = 0.018 \, \text{kg}

  • Initial velocity of the bullet, vi=240.00m/sv_i = 240.00 \, \text{m/s}

  • Final velocity of the bullet, vf=0m/sv_f = 0 \, \text{m/s} (since it comes to rest)

  • Penetration distance, d=35.00cm=0.35md = 35.00 \, \text{cm} = 0.35 \, \text{m}

Solution:

  1. Calculate the initial kinetic energy (KE) of the bullet:

    KE=12mvi2\text{KE} = \frac{1}{2} m v_i^2KE=12×0.018kg×(240.00m/s)2\text{KE} = \frac{1}{2} \times 0.018 \, \text{kg} \times (240.00 \, \text{m/s})^2KE=0.009×57600\text{KE} = 0.009 \times 57600KE=518.4J\text{KE} = 518.4 \, \text{J}

  2. Apply the work-energy principle:

    The work done by the friction force (FfrictionF_{\text{friction}}) brings the bullet to rest. This work is equal to the initial kinetic energy of the bullet.

    Work=Ffriction×d\text{Work} = F_{\text{friction}} \times d

    Since the work done by friction is equal to the initial kinetic energy:

    Ffriction×d=KEF_{\text{friction}} \times d = \text{KE}

    Solving for FfrictionF_{\text{friction}}:

    Ffriction=KEdF_{\text{friction}} = \frac{\text{KE}}{d}Ffriction=518.4J0.35mF_{\text{friction}} = \frac{518.4 \, \text{J}}{0.35 \, \text{m}}Ffriction1481.14NF_{\text{friction}} \approx 1481.14 \, \text{N}

Conclusion:

The magnitude of the constant friction force acting on the bullet as it penetrates the sandbag is approximately 1481.14 N.


answered by: Shital Garg